OpenStudy (setsuna-yuregeshi):

I feel stupid for asking a simple question like this: How do you divide a smaller number by a larger one? Like 7/20 or 5/6 or 10/21?

OpenStudy (nir1232):

you put in the calculator has 7/20

OpenStudy (nir1232):

If u want it has a decimal you out a dot(.) behind it

OpenStudy (dumbcow):

this is long division.... _2.85___ 7| 20 -14 ------- 60 -56 ------ 40 -35 ------ 50 and so on ......

OpenStudy (setsuna-yuregeshi):

I mean by hand and no calculator

OpenStudy (magentama202):

@Setsuna-Yuregeshi any question is ok here people just want to help! At least most :) I can try and help in one sec if you still need it.

OpenStudy (setsuna-yuregeshi):

@MAGENTAMA202 I want and need your help desperately

OpenStudy (magentama202):

one sec now...

OpenStudy (magentama202):

Long division is an algorithm that repeats the basic steps of 1) Divide; 2) Multiply; 3) Subtract; 4) Drop down the next digit.

OpenStudy (magentama202):

do you want help with those division equations in particular? @Setsuna-Yuregeshi

OpenStudy (magentama202):

this link may be helpful and if you're a visual learner it has a video as well at the end. Hope this helps! http://www.homeschoolmath.net/teaching/md/how_teach_long_division.php

OpenStudy (magentama202):

@Setsuna-Yuregeshi sorry I have to leave now.

OpenStudy (legomyego180):

Setsuna do you still need help?

OpenStudy (sshayer):



OpenStudy (unklerhaukus):

When you have a big number divided by a small number, it can be simplified into a mixed numeral. e.g. 20 / 7 = ( 2 x 7 + 6 ) / 7 = 2 + 6/7 . This has the form of whole number (the 2), and a fractional remainder term (the + 6/7). As a mixed numeral, the fraction part will now be a small number divided by a large number. When you have a small number divided by a large number (without common factors) the fraction is often the simplest form for the number it represents. Another representation of such a number is as a decimal. For example the expression 7/20 can be converted into decimal notation like this: 7 / 20 = ( 5 x 7) / ( 5 x 20 ) = 35 / 100 = 35 x 1 / 100 = 35 x 0.01 = 0.35 The fractions ( 5/6 or 10/21 ) are perhaps in their simplest form already. (As they repeat infinity in decimal form.)

OpenStudy (setsuna-yuregeshi):

@legomyego Yes, I'd still like help. I don't really get how to divide a number by a larger number